# Tensor-Var: Variational Data Assimilation in Tensor Product Feature Space

### v2

@@ Abstract
Variational data assimilation estimates the dynamical system states by minimizing a cost function that fits the numerical models with observational data. The widely used method, four-dimensional variational assimilation (4D-Var), has two primary challenges: (1) computationally demanding for complex nonlinear systems and (2) relying on state-observation mappings, which are often not perfectly known. Deep learning (DL) has been used as a more expressive class of efficient model approximators to address these challenges. However, integrating such models into 4D-Var remains challenging due to their inherent nonlinearities and the lack of theoretical guarantees for consistency in assimilation results. In this paper, we propose Tensor-Var to address these challenges using kernel Conditional Mean Embedding (CME). Tensor-Var improves optimization efficiency by characterizing system dynamics and state-observation mappings as linear operators, leading to a convex cost function in the feature space. Furthermore, our method provides a new perspective to incorporate CME into 4D-Var, offering theoretical guarantees of consistent assimilation results between the original and feature spaces. To improve scalability, we propose a method to learn deep features (DFs) using neural networks within the Tensor-Var framework. Experiments on chaotic systems and global weather prediction with real-time observations show that Tensor-Var outperforms conventional and DL hybrid 4D-Var baselines in accuracy while achieving efficiency comparable to the static 3D-Var method. Our implementation is available at: https://github.com/yyimingucl/TensorVar.
@@ 1 Introduction
Forecasting of dynamical systems is an initial value problem of practical significance. Many real-world systems, such as the ocean and atmosphere, are chaotic, which means minor errors of current estimations in computational models can lead to rapid divergence and substantial forecasting errors. In this regard, data assimilation (DA) (Law et al., 2015; Asch et al., 2016) uses observation data to continuously calibrate models, improving forecast accuracy.
@@
Various DA methods have been proposed to deal with different types of observation data and system dynamics. Among these methods, 4D variational (4D-Var) data assimilation has been considered cutting-edge and effectively used in real-world applications like numerical weather prediction (NWP) systems (Browne et al., 2019; Milan et al., 2020). The 4D-Var minimizes a quadratic cost function that finds the optimal match between system states and observations (Asch et al., 2016). While effective in NWP, there are two critical limitations for their applications: (1) Numerical models for complex, nonlinear systems are often inefficient for real-time assimilation and forecasting. (2) Observations are often noisy, incomplete representations of the states, even without a known state-observation mapping, posing challenges in utilizing the observation. Efforts have integrated DL models to learn an observation (or inverse) model (Frerix et al., 2021; Wang et al., 2022; Liang et al., 2023), addressing the imperfect knowledge of observation models in 4D-Var. While these approaches improve observation utilization, they remain constrained by the complexities of numerical models and learned observation mappings, whereas our approach simplifies them by finding their linear representations. To improve computation efficiency, the state-of-the-art DL models (Vaswani et al., 2017; Chen et al., 2018; Li et al., 2020; Kovachki et al., 2023; Bocquet et al., 2024) are capable of constructing highly nonlinear mappings to surrogate dynamical systems and achieve notable successes in NWP (Bi et al., 2022; Lam et al., 2022; Kurth et al., 2023; Chen et al., 2023; Conti, 2024; Vaughan et al., 2024). However, integrating such models into optimization-based tasks, such as 4D-Var, remains challenging due to their inherent nonlinearities (Janner et al., 2021; Bocquet, 2023). Using auto-differentiation (AD) of DL models in 4D-Var can reduce computational costs and has shown success in simple examples (Geer, 2021; Dong et al., 2022; Cheng et al., 2024); however, the accuracy of AD-derived derivatives remains a concern and its complexity grows with system dimensions (Baydin et al., 2018). Recently, Xiao et al. (2024) applied the AD of a pre-trained weather forecasting model into 4D-Var, forming a self-contained DA framework for Global NWP. However, this approach relies on the well-designed pre-trained model for the forward models and may not generalize easily to other domains. Latent data assimilation (Peyron et al., 2021; Fablet et al., 2021; Melinc & Zaplotnik, 2023; Cheng et al., 2023; Fablet et al., 2023) addresses these challenges by performing DA in a learned low-dimensional latent space. While efficient, these approaches lack theoretical guarantees for the consistency of 4D-Var solutions between the latent and original spaces.
@@
In this paper, we introduce Tensor-Var, a framework for learning linear representation of the DA systems by using kernel conditional mean embedding (CME). Unlike existing DL-based DA methods face challenges with nonlinearity and non-convexity, we propose a new perspective from CME to linearize nonlinear dynamics, resulting in a convex cost function in the feature space. To best of our knowledge, our work is first attempt to integrate CME into 4D-Var for linear representation with convex cost function, greatly improving optimization efficiency and convergence. To address the challenges from incomplete observations, we derive an inverse observation operator that incorporates histories to infer the system state, thereby improving accuracy and robustness. Moreover, we provide a theoretical analysis that demonstrates the existence of a linear representation of the system under kernel features and the consistency of 4D-Var solution across original and latent spaces. A key challenge in extending CME to practical variational DA is scalability. To overcome this, our approach learns adaptive deep features (DFs) that map data into a fixed-dimensional feature space, reducing the computational complexities. Our experiments on two chaotic systems and two global NWP applications demonstrate that Tensor-Var outperforms conventional and ML-hybrid variational DA baselines, including operational and DL-hybrid methods, in accuracy and computational efficiency, showing the advantages of linearizing the DA systems through Tensor-Var.
2 Background and Problem Formulation

### v3

# Tensor-Var: Efficient Four-Dimensional Variational Data Assimilation

@@ Abstract
Variational data assimilation estimates the dynamical system states by minimizing a cost function that fits the numerical models with the observational data. Although four-dimensional variational assimilation (4D-Var) is widely used, it faces high computational costs in complex nonlinear systems and depends on imperfect state-observation mappings. Deep learning (DL) offers more expressive approximators, while integrating DL models into 4D-Var is challenging due to their nonlinearities and lack of theoretical guarantees in assimilation results. In this paper, we propose Tensor-Var, a novel framework that integrates kernel conditional mean embedding (CME) with 4D-Var to linearize nonlinear dynamics, achieving convex optimization in a learned feature space. Moreover, our method provides a new perspective for solving 4D-Var in a linear way, offering theoretical guarantees of consistent assimilation results between the original and feature spaces. To handle large-scale problems, we propose a method to learn deep features (DFs) using neural networks within the Tensor-Var framework. Experiments on chaotic systems and global weather prediction with real-time observations show that Tensor-Var outperforms conventional and DL hybrid 4D-Var baselines in accuracy while achieving a 10- to 20-fold speed improvement.
Machine Learning, ICML
@@ 1 Introduction
Forecasting of dynamical systems is an initial value problem of practical significance. Many real-world systems, such as the ocean and atmosphere, are chaotic, which means that minor errors in current estimations in computational models can lead to rapid divergence and substantial forecast errors (Coveney, 2024). In this regard, data assimilation (DA) (Law et al., 2015; Asch et al., 2016) uses observation data to continuously calibrate models, improving forecast accuracy.
Figure 1: Demonstration of Tensor-Var: A DA system with nonlinear dynamical model, observation model, and non-convex cost function (bottom) can be represented linearly in feature space, resulting in a convex cost function (top).
Various DA methods have been proposed to deal with different types of observation data and system dynamics. Among these methods, 4D variational (4D-Var) data assimilation has been considered cutting-edge and used effectively in real-world applications such as numerical weather prediction (NWP) (Browne et al., 2019; Milan et al., 2020). 4D-Var minimizes a quadratic cost function that finds the optimal match between system states and observations (Asch et al., 2016). While effective in NWP, there are two critical limitations for their applications: (1) Numerical models for complex, nonlinear systems are often inefficient for real-time assimilation and forecasting. (2) Observations are often noisy, incomplete representations of the states, or even without a known state-observation mapping, posing challenges in utilizing the observations.
@@
Efforts have been made to integrate DL models to learn an observation (or inverse) model (Frerix et al., 2021; Wang et al., 2022; Liang et al., 2023), addressing the imperfect knowledge of observation models in 4D-Var. While these approaches improve observation utilization, they remain constrained by the complexities of numerical models and learned observation mappings, whereas our approach simplifies them by finding their linear representations. To improve computation efficiency, state-of-the-art DL models (Vaswani et al., 2017; Chen et al., 2018; Li et al., 2020; Kovachki et al., 2023; Cheng et al., 2025; He et al., 2025) are capable of constructing highly nonlinear mappings to surrogate dynamical systems and achieve notable successes in NWP (Bi et al., 2022; Lam et al., 2022; Kurth et al., 2023; Chen et al., 2023; Conti, 2024; Vaughan et al., 2024). However, integrating such models into optimization-based tasks, such as 4D-Var, remains challenging due to their inherent non-linearities (Janner et al., 2021; Bocquet, 2023; Bocquet et al., 2024). Using auto-differentiation (AD) of DL models in 4D-Var can reduce computational costs and has shown success in simple examples (Geer, 2021; Dong et al., 2022; Cheng et al., 2024). However, concerns remain regarding the accuracy of AD-derived derivatives, and their computational complexity increases with the model scale (Baydin et al., 2018). Recently, (Xiao et al., 2024) applied AD of a pre-trained weather forecasting model in 4D-Var, forming a self-contained DA framework for Global NWP. This approach demonstrates promising performance in large-scale settings, but relies on a well-designed pre-trained model and may not generalize easily to other domains. Latent DA (Peyron et al., 2021; Fablet et al., 2021; Melinc & Zaplotnik, 2023; Cheng et al., 2023a; Fablet et al., 2023) addresses these challenges by performing DA in a learned low-dimensional latent space. Although efficient, these approaches lack theoretical guarantees for the consistency of 4D-Var solutions between the latent and original spaces.
@@
In this paper, we introduce Tensor-Var, a framework that linearizes nonlinear dynamics via kernel embedding, enabling convex optimization in feature space. Our approach addresses efficiency and non-convex challenges in existing variational DA methods. To our knowledge, Tensor-Var is the first attempt to identify the linear feature space for DA systems with a convex cost function, enabling an efficient solution to the 4D-Var problem. To handle incomplete observations, we derive an inverse observation operator that incorporates historical observations to infer the system states. Moreover, we provide a theoretical analysis that demonstrates the existence of a linear representation of the system under kernel features and the consistency of the 4D-Var solution across original and feature spaces. A key challenge in extending the kernel embedding to practical variational DA is scalability. To overcome this, our approach learns adaptive deep features (DFs) that map data into a fixed-dimensional feature space, reducing computational complexities. Our experiments on two chaotic systems and two global NWP applications show that Tensor-Var outperforms both conventional and ML-hybrid variational DA baselines in accuracy and computational efficiency, demonstrating the advantages of linearizing DA systems.
2 Preliminary
Notation. Let $S$ and $O$ be random variables representing the state and observation, respectively, with their realizations $s$ and $o$ taking values in the spaces $\mathcal{S}\subset\mathbb{R}^{n_{s}}$ and $\mathcal{O}\subset\mathbb{R}^{n_{o}}$, where $n_{s}$ and $n_{o}$ denote the dimensions of these spaces. The distribution of $S$ is denoted by $\mathbb{P}_{S}$, the joint distribution over $(S,O)$ by $\mathbb{P}_{SO}$, and the conditional distribution of $S$ given $O$ by $\mathbb{P}_{S|O}$. A sequence of states over time steps from $1$ to $t$ is denoted by $s_{1:t}=(s_{1},\ldots,s_{t})$.
2.1 4D variational data assimilation
Consider a dynamical system in discrete-time comprising a dynamical model and observation model:
in which $F$ is the dynamical model that advances the state $s_{t}$ to $s_{t+1}$, and $G$ is the observation model that maps the $s_{t}$ to the observation $o_{t}$. The noise components $\epsilon^{s}_{t},\epsilon^{o}_{t}$ such that $\epsilon^{s}_{t}\sim\mathcal{N}(0,Q)$ and $\epsilon^{o}_{t}\sim\mathcal{N}(0,R)$. The objective of 4D-variational DA is to minimize a cost function:
$\begin{split}J(s_{0:T})=\|s_{0}-s^{b}_{0}\|^{2}_{B^{-1}}&+\sum_{t=0}^{T}\|o_{t}-G(s_{t})\|^{2}_{R^{-1}}\\ &+\sum_{t=1}^{T}\|s_{t}-F(s_{t-1})\|^{2}_{Q^{-1}},\end{split}$ (2)
in which $s_{0}^{b}$ is a prior guess for the initial state $s_{0}$, with $B$ as the background covariance matrix representing the uncertainty, i.e., $s_{0}\sim\mathcal{N}(s_{0}^{b},B)$. The second and third terms account for errors in the observation and dynamical models in (1). Solving the minimization (2) in a high-dimensional nonlinear dynamical system with incomplete observations can be challenging. In this paper, we aim to characterize the nonlinear dynamical and observation models in (1) using linear models, allowing efficient optimization and forecasting. In this sense, we use the CME that transforms the nonlinear system (2) into a linear system in a feature space.
2.2 Kernel conditional mean embedding
A Reproducing Kernel Hilbert Space (RKHS) $\mathbb{H}_{S}$ on $\mathcal{S}$ with kernel $k_{S}$ is a Hilbert space, satisfying the reproducing property (Schölkopf & Smola, 2002). Let $k_{S}$ and $k_{O}$ be positive-definite (pd) kernels on $\mathcal{S}$ and $\mathcal{O}$ with RKHS $\mathbb{H}_{S}$ and $\mathbb{H}_{O}$. We denote the kernel-induced features as $\phi_{O}(o_{t})=k_{O}(o_{t},\cdot)$ and $\phi_{S}(s_{t})=k_{S}(s_{t},\cdot)$, referring the $\mathbb{H}_{S}$ and $\mathbb{H}_{O}$ as feature spaces respectively.
Kernel mean embedding of distribution $\mathbb{P}_{S}$ is defined as the expectation of the feature $\phi_{S}(S):\mu_{\mathbb{P}_{S}}=\mathbb{E}_{S}[\phi_{S}(S)]\in\mathbb{H}_{S}$ and always exists for bounded kernels (Fukumizu et al., 2011). With the reproducing property of RKHS, expectations of the function $f$ can easily be computed as $\mathbb{E}_{S}(f(S))=\langle\mu_{\mathbb{P}_{S}},f\rangle_{\mathbb{H}_{S}}$ for any $f\in\mathbb{H}_{S}$. For characteristic kernels, these embeddings are injective, uniquely determining the probability distribution (Berlinet & Thomas-Agnan, 2011). In addition to mean embedding, we will need the (uncentered) cross-covariance operator (Baker, 1973) $\mathcal{C}_{SO}\colon\mathbb{H}_{O}\rightarrow\mathbb{H}_{S}$, defined as $\mathcal{C}_{SO}=\mathbb{E}_{SO}[\phi_{S}(S)\otimes\phi_{O}(O)]$, where $\otimes$ denotes the tensor product. This operator always exists for bounded kernels and can also be viewed as the embedding of the joint distribution $\mathbb{P}_{SO}$, represented as an element in the tensor product space $\mathbb{H}_{S}\otimes\mathbb{H}_{O}$. Similarly, the covariance operator $\mathcal{C}_{SS}\colon\mathbb{H}_{S}\to\mathbb{H}_{S}$ is defined as $\mathcal{C}_{SS}=\mathbb{E}\big{[}\phi_{S}(s)\otimes\phi_{S}(s)\big{]}$. These operators extend the concepts of covariance matrices from finite-dimensional spaces to infinite kernel feature spaces.
Conditional mean embedding.
To represent the dynamical and observation models that arise in (1), the conditional mean embedding (CME) plays an important role. For a conditional distribution $\mathbb{P}_{S|o}$ of $S$. The CME of $\mathbb{P}_{S|O}$ is defined as
$\mu_{\mathbb{P}_{S|o}}=\mathbb{E}_{S|o}[\phi_{S}(S)|O=o]\quad\text{for any }o\in\mathcal{O},$ (3)
requiring an operator $\mathcal{C}_{S|O}\colon\mathbb{H}_{O}\rightarrow\mathbb{H}_{S}$ such that (1) $\mu_{\mathbb{P}_{S|o}}=\mathcal{C}_{S|O}\phi_{O}(o)$, and (2) $\langle f,\mu_{\mathbb{P}_{S|o}}\rangle_{\mathbb{H}_{S}}=\mathbb{E}_{S|o}[f(S)|O=o]$ for any $f\in\mathbb{H}_{S}$. Under standard assumptions111(1) $\mathcal{C}_{SS}$ is injective, and (2) $\mathbb{E}_{S|o}[f(S)|O=o]\in\mathbb{H}_{S}$ for any $f\in\mathbb{H}_{S}$ and $o\in\mathcal{O}$., the CME operator can be expressed as $\mathcal{C}_{S|O}=\mathcal{C}_{SO}\mathcal{C}_{OO}^{-1}$ (Fukumizu et al., 2004; Song et al., 2009). Given i.i.d samples $\{(s_{i},o_{i})\}_{i=1}^{n}\sim\mathbb{P}_{SO}$, an empirical estimate of the CME operator can be obtained as
$\hat{\mathcal{C}}_{S|O}=\hat{\mathcal{C}}_{SO}(\hat{\mathcal{C}}_{OO}+\lambda I)^{-1}=\Phi_{S}(K_{O}+\lambda I)^{-1}\Phi_{O}^{T},$ (4)
where $\lambda$ is the regularization parameter, $K_{O}$ is the Gram matrix $[K_{O}]_{ij}=k_{O}(o_{i},o_{j})$, and $\Phi_{S}=[\phi_{S}(s_{1}),...,\phi_{S}(s_{n})]$ and $\Phi_{O}=[\phi_{O}(o_{1}),...,\phi_{O}(o_{n})]$ are the feature matrices stacked by columns.
Table 1: Comparison of Tensor-Var with other algorithms. Here, $\epsilon$ is the error threshold and $T$ is the length of the assimilation window.
Space Transformation Non-Space Transformation
Methods Tensor-Var Latent 4D-Var 4D-Var (w/o or w/ Adjoint) Frerix et al. (2021)
Convexity
Convergence Rate Linear Sublinear Sublinear Sublinear
Iteration Time $\mathcal{O}\left(\log\left(1/\epsilon\right)\right)$ $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}\left(1/\epsilon^{2}\right)$ $\mathcal{O}\left(1/\epsilon^{2}\right)$
Computational Complexity $\mathcal{O}\left(Td_{s}\log\left(1/\epsilon\right)\right)$ $\mathcal{O}\left((Td_{s})^{2}/\epsilon^{2}\right)$ w/o Adjoint: $\mathcal{O}\left((Tn_{s})^{2}/\epsilon^{2}\right)$
w/ Adjoint: $\mathcal{O}\left(Tn_{s}/\epsilon^{2}\right)$
$\mathcal{O}\left((Tn_{s})^{2}/\epsilon^{2}\right)$
3 Method
In this section, we introduce our Tensor-Var approach, which embeds 4D-Var into the kernel feature space and provides a theoretical analysis demonstrating the existence of linear dynamics with consistent convergence between the original and feature space solutions. To effectively address incomplete observations, we propose an inverse observation operator that leverages consecutive historical observations. Additionally, we propose a method to learn adaptive deep features (DFs) using neural networks within the Tensor-Var framework, improving real-world applicability. Finally, we analyze optimization performance of Tensor-Var compared to existing 4D-Var based methods, demonstrating its efficiency.
3.1 CME of 4D-Var in RKHS
Having introduced the necessary tools for manipulating kernel embeddings, we now focus on learning the linearized models of the system in (1).
CME of dynamical model.
Let $S^{+}$ be the one-step forward of $S$. Given the dynamical model $F$ in (1) and the kernel feature $\phi_{S}$, the CME operator $\mathcal{C}_{S^{+}|S}$ can be recognized as the best linear approximation in the feature space $\mathbb{H}_{S}$ that minimizes the regression residual $\mathbb{E}\big{[}\|\phi_{S}(S^{+})-\mathcal{C}\phi_{S}(S)\|^{2}\big{]}$. Given finite data $\{(s_{i}^{+},s_{i})\}_{i=1}^{N}$ sliced from the system trajectory $s_{1:N+1}$, we can obtain the empirical estimate $\hat{\mathcal{C}}_{S^{+}|S}$ as (4) with theoretical support of convergence (Fukumizu et al., 2013; Klus et al., 2020). The CME operator $\hat{\mathcal{C}}_{S^{+}|S}$ effectively characterizes the system dynamics as a linear model and simplifies the 4D-Var as a convex optimization in the feature space $\mathbb{H}_{S}$.
CME of inverse observation model.
Analogous to the dynamical model, the observation model $G$ in (1) can be linearized by the CME operator $\mathcal{C}_{O|S}$, which has been used as observation models for filtering algorithms (Song et al., 2009; Fukumizu et al., 2013; Kanagawa et al., 2016; Gebhardt et al., 2019). Most approaches assume a complete observation setting, where the observations can fully determine the state. In practice, observations are often incomplete representations of states, with $n_{s}>n_{o}$, leading to underdetermined systems (Liu et al., 2022). In such systems, the lack of a bijective mapping between the state and observation spaces means that observations cannot uniquely determine the system’s state. As a result, the optimization problem in 4D-Var may produce suboptimal solutions (Asch et al., 2016). It is theoretically challenging to distinguish any two mixtures of states based on a single-step observation if $n_{s}>n_{o}$. By introducing the past $m$ consecutive observations as history $h_{t}=o_{t-m-1:t-1}\in\mathbb{R}^{m\times n_{o}}$, the joint information from history and current observation is enough to estimate the system state. The choice of history length is critical: Too short lacks sufficient information, while too long is inefficient. Empirically, we performed an ablation study to assess the effects of history length, as detailed in the Subsection 4.4.
To effectively incorporate history, we introduce another kernel feature $\phi_{H}$ with an induced RKHS $\mathbb{H}_{H}$. The space $\mathbb{H}_{OH}=\mathbb{H}_{O}\otimes\mathbb{H}_{H}$ is called a tensor product RKHS on $O\times H$ with associated kernel feature $\phi_{OH}(o,h)=\phi_{O}(o)\otimes\phi_{H}(h)$. As shown in (Song et al., 2013; Muandet et al., 2017), the CME operator can be extended to high-order features, allowing us to embed the joint distributions over $s_{t}$, $o_{t}$, and $h_{t}$ into the feature space (Song et al., 2009, 2013). The tensor product feature $\phi_{OH}$ captures the high-order dependencies between observations and the history. The CME operator $\mathcal{C}_{S|OH}$ is the linear inverse observation model that minimizes the state estimation error given observations and history. Given dataset $\{(s_{i},o_{i},h_{i})\}_{i=1}^{N}$, the empirical counterpart $\hat{\mathcal{C}}_{S|OH}=\hat{\mathcal{C}}_{SOH}(\hat{\mathcal{C}}_{(OH)(OH)}+\lambda I)^{-1}$ follows the same way as (4). The $\hat{\mathcal{C}}_{SOH}$ is the empirical high order tensor, e.g., $\hat{\mathcal{C}}_{SOH}=\sum_{i=1}^{N}\phi_{S}(s_{i})\otimes\phi_{O}(o_{i})\otimes\phi_{H}(h_{i})$.
Feature space 4D-Var.
Using the kernel features, we linearize the original nonlinear dynamics and observations in the feature space $\mathbb{H}_{S}$. This transformation enables us to reformulate the 4D-Var optimization objective (2) into the feature space and optimize over a sequence of elements $z_{0:T}$ in feature space $\mathbb{H}_{S}$:
$\begin{split}\min_{z_{0:T}}\,\,&\|z_{0}-\phi_{S}(s_{0}^{b})\|^{2}_{\mathcal{B}^{-1}}+\sum_{t=0}^{T-1}\|z_{t+1}-\mathcal{C}_{S^{+}|S}z_{t}\|^{2}_{\mathcal{Q}^{-1}}\\ &+\sum_{t=0}^{T}\|z_{t}-\mathcal{C}_{S|OH}\phi_{OH}(o_{t},h_{t})\|^{2}_{\mathcal{R}^{-1}},\end{split}$ (5)
where the $\mathcal{B}$, $\mathcal{R}$, and $\mathcal{Q}$ are the covariance operators for the background error, observation error, and model error in the feature space. In this work, we estimate the three operators as the empirical error covariance matrices from the training dataset (the explicit estimation can be found in Appendix D.1). We present the pseudo-algorithms in Appendix D.2, as shown in Algorithms 2 and 3.
3.2 Learning the deep features within Tensor-Var
Using the pre-determined kernel features has theoretical guarantees. However, it maps data into an infinite-dimensional feature space, e.g., radial basis function, making estimation challenging due to polynomial scaling with sample size. On the other hand, these feature maps struggle with irregular or high-dimensional data, often resulting in poor performance. Learned deep features (DFs) have emerged as alternatives to generic kernel features (Xu et al., 2022; Kostic et al., 2023; Shimizu et al., 2024), projecting the data into a fixed-dimensional feature space. To improve scalability, we integrate DFs with the Tensor-Var framework and validate their effectiveness through experiments.
Learning the state feature.
Recall that the CME operator $\mathcal{C}_{S^{+}|S}$ is the best linear approximation of system dynamics in feature space. We propose to jointly learn the feature $\phi_{\theta_{S}}\colon\mathbb{R}^{n_{s}}\rightarrow\mathbb{R}^{d_{s}}$ with $\hat{\mathcal{\mathcal{C}}}_{S^{+}|S}$ by minimizing the loss $L(\theta_{S})=\min_{\mathcal{C}\in\mathbb{R}^{d_{s}\times d_{s}}}L(\mathcal{C},\theta_{S})=\mathbb{E}\big{[}\|\phi_{\theta_{S}}(s^{+})-\hat{\mathcal{C}}_{S^{+}|S}\phi_{\theta_{S}}(s)\|^{2}\big{]}$. Predictions using $\hat{\mathcal{C}}_{S^{+}|S}$ are in the feature space; however, for DA problems, reconstruction to the original state space is required, which is known as the preimage problem (Honeine & Richard, 2011). Here, we learn an inverse feature $\phi^{\dagger}_{\theta^{\prime}_{S}}$ to solve the preimage problem during training, avoiding repeated optimization whenever computing preimages. The final training loss is the combination of the two terms as
$\begin{split}L(\theta_{S},\theta_{S}^{\prime})&=\mathbb{E}\big{[}\|\phi_{\theta_{S}}(s^{+})-\hat{\mathcal{C}}_{S^{+}|S}\phi_{\theta_{S}}(s)\|^{2}\big{]}\\ &+w\,\mathbb{E}\Big{[}\|s-\phi^{\dagger}_{\theta^{\prime}_{S}}\big{(}\phi_{\theta_{S}}(s)\big{)}\|^{2}\Big{]},\end{split}$
where $w\in(0,1]$ is a weighting coefficient, and $\hat{\mathcal{C}}_{S^{+}|S}$ is computed as the CME over training batches. Note that using DFs corresponds to a linear kernel in the learned feature space, where $k_{\theta_{S}}(s_{i},s_{j})=\phi_{\theta_{S}}(s_{i})^{T}\phi_{\theta_{S}}(s_{j})$.
Learning the observation and history features.
Similar to learning the state feature, $\mathcal{C}_{S|OH}$ is the minimizer of the regression problem mapping the tensor product of observation and history features to the state feature. In this phase, we learn the DFs for observation $\phi_{\theta_{O}}\colon\mathbb{R}^{n_{o}}\rightarrow\mathbb{R}^{d_{o}}$, history $\phi_{\theta_{H}}\colon\mathbb{R}^{n_{h}}\rightarrow\mathbb{R}^{d_{h}}$, and $\mathcal{C}_{S|OH}$ jointly with the loss function:
where $\hat{\mathcal{C}}_{S|OH}$ is computed over training batches in parallel, with the $\otimes$ denoting the Kronecker product in practice.
Improved 4D-Var Optimization.
Tensor-Var improves 4D-Var optimization by transforming the nonlinear problem into a linear convex optimization, enabling global optimality and faster convergence in feature space. Unlike conventional 4D-Var, which relies on costly Hessian computations and suffers from sublinear convergence, Tensor-Var achieves linear convergence through linearized dynamics $z_{t+1}=\mathcal{C}_{S^{+}|S}z_{t}$. Table 1 presents a theoretical comparison of optimization properties between Tensor-Var and existing 4D-Var-based methods, including ML-hybrid 4D-Var (Frerix et al., 2021), latent 4D-Var (Cheng et al., 2024), and traditional 4D-Var with or without an adjoint model. As shown in Table 1, due to its linear dynamics, its iteration time scales as $\mathcal{O}(\log(1/\epsilon))$ with error threshold $\epsilon$, which is more efficient than $\mathcal{O}(1/\epsilon^{2})$ the complexity of most 4D-Var algorithms, as supported by empirical results in Section 4.1. Notably, there is an additional error source from the feature mapping, which is not explicitly discussed here; a more in-depth analysis can be found in Appendix B.
3.3 Theoretical analysis.
In Section 3.1, we discuss how the dynamical system $F$ in (1) can be embedded as a linear system in the feature space. However, two important questions remain: 1) Does such a linear dynamical system exist? 2) Are the solutions of the original and feature space 4D-Var consistent? In this section, we provide affirmative answers to both questions using the theory of Kazantzis-Kravaris/Luenberger (KKL) observers (Andrieu & Praly, 2006). We give a road map of the theoretical analysis with the main results and refer the reader to Appendix C for details.
Under the mild assumptions that (1) the dynamical model $F$ is first-order differentiable and (2) the kernel features are all first-order differentiable, the kernel feature $\phi_{S}$ satisfies the necessary conditions as the state transformation in the KKL observer framework. This transformation enables us to represent the nonlinear dynamical system as a linear system in a higher-dimensional feature space, as established by KKL observer theory (Tran & Bernard, 2023). This result confirms the existence of such a linear system, thus answering question 1). The KKL observer theory provides a theoretical foundation for our approach, bridging nonlinear dynamics and linear 4D-Var methods (Andrieu & Praly, 2006). A detailed derivation proving that $\phi_{S}$ satisfies the conditions as the state transformation of the KKL observer can be found in Appendix C.
We consider the nonlinear system in (1) within a compact state space and assume that the cost function in 2 has a unique solution. Given that $\phi_{S}$ is a state transformation in the KKL observer, the system in the original state space can be represented linearly in the feature space. The solution in the feature space has a consistent convergence to the unique solution of the original 4D-Var problem, minimizing with respect to the cost function 5, answering question 2). A formal theorem with detailed proofs can be found in Theorem C.9 in Appendix C.2.
4 Experiment
Table 2: Comparison of DA performances. All baseline methods use the strong-constraint 4D-Var objective, while our approach uses the weak-constraint 4D-Var objective. Evaluation times and iterations are reported as each assimilation window’s mean and standard deviation. Our method consistently outperforms the baselines across all benchmark domains.